OpenStudy (anonymous):

Diana invested $3000 in a savings account for 3 years. She earned $450 in interest over that time period. What interest rate did she earn? Use the formula I=Prt to find your answer, where I is interest, P is principal, r is rate and t is time. Enter your solution in decimal form rounded to the nearest hundredth. For example, if your solution is 12%, you would enter 0.12.

OpenStudy (anonymous):

3000*X*3=450 9000*X=450 X=\[\frac{ 450 }{ 9000 }\] X=.05 Her interest rate for 3 years was 5%.

OpenStudy (tkhunny):

What is "X"? I = Prt We are given: "Diana invested $3000 in a savings account" P = $3000 "for 3 years." t = 3 "She earned $450 in interest over that time period." I = $450 Thus: $450 = $3000 * r * 3 Solve for r.

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

X is suppose to be r, my fault for the confusion.

OpenStudy (anonymous):

thats 0.05 in a decimal ?

OpenStudy (anonymous):

yes. .05 as a percent is 5%. But as a decimal is 0.05